OpenStudy (anonymous):

For the equation y=(x+27)^2/3 -1, find the coordinates of the point(s) at which the graph crosses or touches the coordinate axes, if they exist.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x=-26

OpenStudy (anonymous):

(-26,0) and (0,8)

OpenStudy (anonymous):

The answers it gives me are (-26,0),(-28,0),(0,8)

OpenStudy (anonymous):

yes thats true...

OpenStudy (anonymous):

however i have no clue how to get to those answers myself

OpenStudy (anonymous):

put x=0

OpenStudy (anonymous):

y=(27)^3/2-1 y=9-1=8

OpenStudy (anonymous):

yea i got that one

OpenStudy (anonymous):

its the x values that get me

OpenStudy (anonymous):

and put y=0 then 0=(x+27)^2/3-1 1=(x+27)^2/3 1=(x+27)^2 (x+27)=+-1 x=-26 or -28

OpenStudy (anonymous):

i think its clear now...

OpenStudy (anonymous):

thanks alot ill try a few with your examples

OpenStudy (anonymous):

still trying to figure out where the 2/3 plays in lol